OpenStudy (anonymous):

(12/100)87600=x whats x?

OpenStudy (anonymous):

= 12 * 876 = 10512

OpenStudy (anonymous):

12 x 87600 = 12 * 876 = 10512 -- 100

OpenStudy (anonymous):

are you sure this time

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

niceeee

OpenStudy (anonymous):

87600 divided by 100 = 876 and i used my calculator to multipy by 12